# Large Language Models for Autonomous Driving (LLM4AD): Concept, Benchmark, Simulation, and Real-Vehicle Experiment

### v1

@@ Abstract
With the broader usage and highly successful development of Large Language Models (LLMs), there has been a growth of interest and demand for applying LLMs to autonomous driving technology. Driven by their natural language understanding and reasoning ability, LLMs have the potential to enhance various aspects of autonomous driving systems, from perception and scene understanding to language interaction and decision-making. In this paper, we first introduce novel concepts and approaches to designing LLMs for autonomous driving (LLM4AD). Then, we propose a comprehensive benchmark for evaluating the instruction-following abilities of LLMs within the autonomous driving domain. Furthermore, we conduct a series of experiments on both simulation and real-world vehicle platforms, thoroughly evaluating the performance and potential of our LLM4AD systems. Our research highlights the significant potential of LLMs to enhance various aspects of autonomous vehicle technology, from perception and scene understanding to language interaction and decision-making.
@@ I-A Background
  • Contextual Understanding and Reasoning: LLMs provide contextual understanding from diverse sources such as traffic laws and accident reports, allowing generated decisions to guarantee safety and follow local regulations.
  • Zero-Shot and Few-Shot Planning: The zero-shot generalization ability enables LLMs to perform tasks they have not been trained on before [kojima_large_2022]. This means LLM-based autonomous systems can handle uncommon situations with minimal or no prior experience, enabling them to confidently navigate through ‘corner cases.’
@@
Despite the numerous advantages above, LLMs still have some limitations and potential issues. The most obvious one is sometimes they are inapplicable to real-time tasks in autonomous driving. Specifically, LLMs typically require seconds to process textual information. This delay can pose significant safety concerns and risks in situations requiring immediate, real-time decision-making. Additionally, addressing hallucinations is also a main challenge in LLM4AD systems. “Hallucination” refers to instances where LLMs generate output that is factually incorrect, nonsensical, or unrelated to the input prompt. As autonomous driving is a safety-critical task, hallucinations can introduce significant safety and reliability concerns and potentially undermine trust in LLM4AD systems. Furthermore, data privacy and security are other potential issues since LLMs collect and process vast amounts of textual data, including potentially sensitive information about their surroundings, passengers, and driving preferences.
@@ I-B Overview of This Article
Our paper presents a comprehensive study on integrating LLMs into autonomous driving systems, from the proposed concept and the creation of an open dataset to the implementation in simulations and real-vehicle experiments. Sec. II introduces the concept and key elements of our proposed LLM4AD framework. Sec. III develops an open benchmark designed to evaluate the performance of LLMs in various autonomous driving tasks, including dataset, evaluation metrics, and baseline model. Sec. IV validates the proposed concepts through extensive simulations, exploring a key perspective: learning from human feedback. LABEL:sec:real-vehicle presents the implementation of our framework on a real-vehicle platform and discusses the results of a series of experiments conducted to validate the system’s driving and personalization performance. LABEL:sec:discussion analyzes the current limitations and outlines promising future research directions for leveraging LLMs in autonomous driving applications. Finally, LABEL:sec:conclusion summarizes the key contributions and findings of this work.
Figure 1: Our proposed framework is an LLM-driven autonomous driving system. Within our framework, LLMs receive inputs including system messages, human instructions, situation descriptions, and the historical memory of the specific user. The outputs consist of two main components: a generated program and reasoned thoughts. The generated program is subsequently executed within the executor, interacting with the environment.
@@
$P$ Generated Program
$P^{\prime}$ Improved Generated Program
@@ II-E Large Language Models
LLMs serve as the core module in our framework, which will receive all the inputs (history memories (interactions) $H$, system message $S$, situation description $C$, human instruction $I$, and human evaluation $F$) and generate the textual outputs (LMPs $P$ and reasoning thoughts $R$). One important note is that we employ the chain-of-thought prompting [wei_chain--thought_2022] technique, which serves as a guiding signal to ensure alignment with human-like reasoning and practical driving considerations. Chain-of-thought presents a structured sequence of reasoning examples to LLMs, bridging existing knowledge gaps. By offering a series of logical and connected steps, LLMs can perform well in complex driving scenarios more effectively.
II-F Generated Program
@@ II-G Output Thought
By employing chain-of-thought prompting [wei_chain--thought_2022], LLMs not only generate the program code but also provide a step-by-step explanation of the thought process used to reach that solution. These chains of thoughts represent LLMs’ reasoning behind each decision, such as “Since the command is ‘hurry,’ I will increase the target velocity” or “To turn left, I need to adjust the steering angle based on the current speed.” The output thoughts $R$ accompany the generated programs $P$, offering insights into how LLMs interpret natural language commands in the context of the driving situation description to produce precise control values like velocities or steering angles. This output thought improves the transparency and interpretability of the LLM4AD system’s decision-making process.
@@ II-H Executor
III An Benchmark Dataset of LLM4AD
In this section, we introduce LaMPilot-Bench, the first benchmark of its kind for evaluating the instruction-following capabilities of LLM-based agents in autonomous driving. LaMPilot-Bench consists of three key components: a simulator, a dataset, and an evaluator.
III-A Simulator
The LaMPilot-Bench simulator is built upon HighwayEnv [leurent_environment_2018], a widely used platform for autonomous driving research and tactical decision-making. HighwayEnv offers various driving models and simulates realistic multi-vehicle interactions. We extend HighwayEnv with interfaces suitable for LLM-based agents and implement custom intersections to diversify the driving scenarios.
III-B Dataset
The LaMPilot dataset consists of 4,900 semi-human-annotated traffic scenes, with a subset of 500 samples split as the test set. Each data sample includes:
  • An instruction $I$: a high-level task description.
  • An initial state: used to initialize the simulator.
  • Goal state criteria: aligned with the instruction $I$.
The dataset covers diverse driving scenarios, as shown in Tab. II. For each driving scenario, LaMPilot includes various situations. Taking turning scenarios as an example, the diversity is reflected in several variables, such as the ego vehicle’s initial position and state, the specific task (turning left/right or going straight), the number of other vehicles, and their positions and states. The driving model parameters for other vehicles are randomly initialized, and each scenario is assigned a random seed. The dataset also contains a variety of instructions reflecting realistic in-cabin human commands categorized by maneuver types (e.g., routing, lane changing, overtaking) and scenario types (highway and intersection). Fig. 2 shows the distribution of the first four words in the instructions, while main statistics are provided in Tab. II.
@@ III-B Dataset
III-C Evaluator
The LaMPilot-Bench evaluator incorporates metrics to assess the safety and efficiency of the agent driving policies. Time-to-collision (TTC) is used to measure the vehicle’s ability to maintain safe distances and avoid collisions. In a scenario with $n+1$ vehicles (ego vehicle labeled as 00), the TTC ($\tau_{i}$) with vehicle $i$ is calculated using the ego vehicle’s longitudinal velocity $\mathbf{v}_{0}$ and position $\mathbf{p}_{0}$, and vehicle $i$’s longitudinal velocity $\mathbf{v}_{i}$ and position $\mathbf{p}_{i}$ ($1\leq i\leq n$):
@@ III-C Evaluator
Speed variance (SV) is considered as another safety metric [garber_factors_1989]. The ego vehicle’s speed standard deviation $\sigma_{0}$ is calculated as:
@@
A TE score closer to 100 indicates a more efficient policy.
A task is considered successfully completed when the agent achieves the objectives specified in the instruction while maintaining safety (i.e., avoiding collisions) and efficiency (i.e., finishing within the allotted time). For example, a lane change task is completed when the vehicle is in the target lane with its heading aligned with the lane’s direction within a specified threshold. The final score aggregates all individual metrics, weighted according to their importance:
@@
III-D Experiment Setup
The experiments are designed to establish baselines for the LaMPilot-Bench. The main objectives are to assess the performance of LLM-based agents in interpreting human instructions within driving contexts and to evaluate the capability of LLMs to generate code for motion planning using provided functional primitives.
We conduct benchmarking using various state-of-the-art large language models, including both open-source and proprietary solutions [^1] OpenAI API accessed in October and November 2023.:
  • Llama 2 [touvron_llama_2023] (llama-2-70b-chat)
  • PaLM 2 [anil_palm_2023] (code-bison)
  • ChatGPT [openai_introducing_2023] (gpt-3.5-turbo)
  • GPT-4 [openai_gpt-4_2023] (gpt-4)
  • GPT-4-Turbo [openai_gpt-4-turbo_2023] (gpt-4-1106-preview)
We set a 60-second time limit for each scenario. If a task is not completed within this time frame, the test case is considered a failure, and the simulation is terminated, resulting in a driving score of 0. For successful cases, the driving score is calculated as follows:
$\operatorname{Driving\ Score}=\frac{\alpha}{N_{s}}\sum_{i=1}^{N_{s}}\operatorname{Score}_{i}-\beta\cdot P_{\operatorname{collision}}$ (9)
Here, $N_{s}$ represents the number of successful test cases, $\alpha$ is the success rate (ranging from 0 to 1), $\beta$ is the collision rate (also between 0 and 1), and $P_{\text{collision}}$ is the penalty factor for collisions (set at 500 in our experiments). The $\text{Score}_{i}$ for each individual sample is calculated based on Eq. 8, with weights $W_{\text{TTC}}$, $W_{\text{SV}}$, and $W_{\text{TE}}$ set to 0.5, 0.3, and 0.2, respectively. The driving score provides a comprehensive assessment of the agent’s performance in LaMPilot-Bench.
III-E Baselines
III-E1 Heuristic Baselines
In our experiments, we employ two rule-based baseline policies: the Intelligent Driver Model (IDM)[treiber_congested_2000] and the Minimizing Overall Braking Induced by Lane Changes (MOBIL) principle[kesting_general_2007]. IDM describes a rule for updating the acceleration of a vehicle to avoid collisions based on the proximity and relative velocity of the vehicle to the object directly in front. MOBIL is an extension of IDM where a lane change is executed if the prospective new lane offers a more favorable driving scenario and the maneuver can be conducted safely. These baselines can be considered as operating on random chance, as the policy is independent of user instructions but follows predefined rules. Additionally, we include a human performance baseline, where a licensed human driver controls the vehicle in the simulation using arrow keys on the keyboard, following the commands and visuals displayed. This baseline provides a reference for human-level performance on LaMPilot-Bench.
III-E2 Zero-Shot and Few-Shot Baselines
We also include the zero-shot and few-shot CoT prompting [wei_chain--thought_2022] as baseline methods. The zero-shot CoT baseline leverages the prompt “How to complete the task step by step.” In the few-shot setting, we follow the standard practice [brown_language_2020] by including 3 human-written code exemplars before the test instance. These in-context examples help the model adapt to the tasks in LaMPilot-Bench. The in-context examples are created by a programmer proficient in Python. They are provided with API descriptions and are allowed to write and test their code using the non-test scenarios. The same set of three examples are used for all test cases. The API Docs are also the same across all test cases for both zero-shot and few-shot settings.
III-E3 Human Feedback Baselines
LLMs have demonstrated proficiency in generating coherent solutions across various tasks without requiring additional fine-tuning. However, when it comes to code generation, especially for complex scenarios, they can produce suboptimal results. The autoregressive nature of these models poses a significant challenge, as tokens generated early in a sequence cannot be modified in the same iteration. This constraint limits the models’ ability to refine initial responses, potentially impacting the effectiveness of the generated code [wang_voyager_2023].
To address these challenges and enhance the performance of LLMs on tasks in LaMPilot-Bench, we include the human-in-the-loop baseline introduced in LABEL:sec:human-in-the-loop-learning, where the LLMs serve not only as planners but also as engines for incorporating human feedback. After an LLM generates a policy program ($P$), it can integrate human feedback ($F$) with context-specific guidance, enabling the LLM to refine its output.
This approach does not involve fine-tuning or training a new LLM. Instead, we leverage the RAG approach [lewis_retrieval-augmented_2020] and introduce a code repository module, which is implemented with the Chroma [chroma_ai-native_2023] vector database. This repository allows for the storage and retrieval of effective code snippets to be used in similar situations. Following Voyager [wang_voyager_2023], we use LLM-generated function descriptions, which are then converted into vectors to serve as database keys. These keys are paired with the corresponding function codes as their values. If the human is satisfied with the policy, the code is committed to the repository for future reference. The human feedback learning process is conducted on the non-test set, involving interaction with 100 random samples for each method to ensure fair comparisons.
III-F Quantitative Results
In this section, we present the experimental results on the LaMPilot-Bench, summarizing the performance of various methods, including heuristic baselines, zero-shot and three-shot baselines, and human feedback baselines. These results are shown in Tab. III.
Rule-Based Methods
@@ III-F Quantitative Results
We first evaluate the performance of rule-based approaches, specifically the IDM and MOBIL algorithms. Both methods achieve a zero collision rate in LaMPilot-Bench. However, it is important to note that these methods operate independently of the provided instructions. Without considering human instructions, IDM and MOBIL achieve success rates of 20.4% and 15.3%, respectively. These results provide a reference point for assessing the effectiveness of LLM-based agents in following human instructions.
LLM-Based Agents
We assess the effectiveness of off-the-shelf LLMs to reason in the context of autonomous driving and following human instructions. In the zero-shot setting, GPT models and PaLM 2, given only the API Docs, driving context, and instructions, achieve an obvious performance advantage compared to rule-based methods in terms of task completion. This improvement can be attributed to the LLMs’ ability to understand and generate code based on natural language instructions, leveraging their extensive pretraining on diverse data sources. However, the zero-shot setting also results in an increase in the collision rate (1%-4%), indicating that the generated code policies do not fully capture the complexities of driving tasks.
When provided with training examples that include exemplar code (three-shot setting), all the evaluated LLMs exhibit notable improvements in completion rates. For instance, Llama 2’s completion rate increases from 20.4% in the zero-shot setting to 63.3% in the three-shot setting. This improvement demonstrates the effectiveness of few-shot learning in adapting LLMs to specific task domains, as the exemplar code provides a clearer understanding of the desired behavior and API usage patterns.

### v3

# Large Language Models for Autonomous Driving (LLM4AD): Concept, Benchmark, Experiments, and Challenges

@@ Abstract
With the broader usage and highly successful development of Large Language Models (LLMs), there has been a growth of interest and demand for applying LLMs to autonomous driving technology. Driven by their natural language understanding and reasoning ability, LLMs have the potential to enhance various aspects of autonomous driving systems, from perception and scene understanding to language interaction and decision-making. In this paper, we first introduce the novel concept of designing LLMs for autonomous driving (LLM4AD). Then, we propose a comprehensive benchmark for evaluating the instruction-following abilities of LLM4AD in simulation. Furthermore, we conduct a series of experiments on real-world vehicle platforms, thoroughly evaluating the performance and potential of our LLM4AD systems. Finally, we envision the main challenges of LLM4AD, including latency, deployment, security and privacy, safety, trust and transparency, and personalization. Our research highlights the significant potential of LLMs to enhance various aspects of autonomous vehicle technology, from perception and scene understanding to language interaction and decision-making.
@@ I Introduction
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=10em,font=,, where level=2text width=11em,font=,, where level=3text width=18em,font=,, [ Large Language Models for Autonomous Driving, ver [ Concept of LLM4AD ] [ Benchmark and Simulation [ LaMPilot-Bench, leaf ] [ CARLA Leaderboard 1.0 Benchmark, leaf ] [ Baselines and Results, leaf ] [ Summary, leaf ] ] [ Real-Vehicle Experiments:
On-Cloud LLM for Personalized
Decision-Making [ Talk2Drive Framework, leaf ] [ Autonomous Driving System, leaf ] [ Experiment Setup, leaf ] [ Experiment Results, leaf ] ] [ Real-Vehicle Experiments:
On-Edge VLM for Personalized
Motion Control [ Problem Statement, leaf ] [ Our Framework, leaf ] [ Real-World Experiment, leaf ] ] [ Challenges [ Latency, leaf ] [ Security and Privacy, leaf ] [ Development, leaf ] [ Safety, leaf ] [ Personalization, leaf ] [ Trust and Transparency, leaf ] ] ]
Figure 1: A condensed overview of this paper.
@@ I-A Background
  • Contextual Understanding and Reasoning: LLMs provide contextual understanding from diverse sources such as traffic rules and accident reports, allowing generated decisions to guarantee safety and follow local regulations.
  • Zero-Shot and Few-Shot Planning: The zero-shot generalization ability enables LLMs to perform tasks they have not been trained on before [kojima_large_2022]. This means LLM-based autonomous systems can handle uncommon situations with minimal or no prior experience, enabling them to confidently navigate through “corner cases.”
@@
Despite the numerous advantages above, LLMs still have some limitations and potential issues. The most obvious one is sometimes they are inapplicable to real-time tasks in autonomous driving. Specifically, LLMs typically require seconds to process textual information. This delay can pose significant safety concerns and risks in situations requiring immediate, real-time decision-making. Additionally, how to avoid the negative effect provided by hallucinations is also a main challenge in LLM4AD systems. “Hallucination” refers to instances where LLMs generate output that is factually incorrect, nonsensical, or unrelated to the input prompt. As autonomous driving is a safety-critical task, hallucinations can introduce significant safety and reliability concerns and potentially undermine trust in LLM4AD systems. Furthermore, data privacy and security are other potential issues since LLMs collect and process vast amounts of textual data, including potentially sensitive information about their surroundings, passengers, and driving preferences.
@@ I-B Overview of This Article
Fig. 1 provides an overview of the reminder structure of this paper. Our paper presents a comprehensive study on integrating LLMs into autonomous driving systems, from the proposed concept and the creation of an open dataset to the implementation in simulations and real-vehicle experiments. Sec. II introduces the concept and key elements of our proposed LLM4AD framework. Sec. III develops an open benchmark designed to evaluate the performance of LLMs in various autonomous driving tasks, which includes a dataset, evaluation metrics, and baseline models. Sec. IV and Sec. V present the implementation of our framework on a real-vehicle platform and discuss the results of a series of experiments conducted to validate the system’s driving and personalization performance. Sec. VI analyzes the current challenges and outlines promising future research directions for leveraging LLMs in autonomous driving applications. Finally, Sec. VII summarizes the key contributions and findings of this work.
Figure 2: The conceptual illustration of the proposed LLM4AD system.
@@
$P$ Generated Program/Policy
$V$ Visual Input
$D$ Scene Description
$A$ Executed Action
@@ II-E Large Language Models
LLMs serve as the core module in our framework, which will receive all the inputs (history interactions) $H$, system message $S$, situation description $C$, human instruction $I$, and human evaluation $F$) and generate the textual outputs (LMPs $P$ and reasoning thoughts $R$). One important note is that we employ the chain-of-thought prompting [wei_chain--thought_2022] technique, which serves as a guiding signal to ensure alignment with human-like reasoning and practical driving considerations.
II-F Generated Policy/Program
@@ II-G Output Thought
By employing chain-of-thought prompting [wei_chain--thought_2022], LLMs not only generate the program code but also provide a step-by-step explanation of the thought process used to reach that solution. These chains of thoughts represent LLMs’ reasoning behind each decision, such as “Since the command is “hurry,” I will increase the target velocity” or “To turn left, I need to adjust the steering angle based on the current speed.” The output thoughts $R$ accompany the generated programs $P$, offering insights into how LLMs interpret natural language commands in the context of the driving situation description to produce precise control values like velocities or steering angles. This output thought improves the transparency and interpretability of the LLM4AD system’s decision-making process.
@@ II-H Executor
III Benchmarking LLM4AD in Simulation
This section introduces two comprehensive benchmarks for evaluating the instruction-following capabilities and performance of LLM-based agents in autonomous driving: the LaMPilot-Bench dataset on the HighwayEnv simulator and the CARLA Leaderboard 1.0 [dosovitskiy_carla_2017].
III-A LaMPilot-Bench
III-A1 Dataset
The LaMPilot-Bench dataset consists of 4.9K semi-human-annotated traffic scenes, with a subset of 500 samples split as the test set. Each data sample includes a high-level task description (instruction), an initial state for initializing the simulator, and goal state criteria aligned with the instruction. The dataset covers diverse driving scenarios, as shown in Tab. II. For each scenario, LaMPilot-Bench includes various situations, such as different ego vehicle initial positions and states, specific tasks (turning left/right or going straight), the number of other vehicles, and their positions and states. The driving model parameters for other vehicles are randomly initialized, and each scenario is assigned a random seed.
@@ III-A1 Dataset
The LaMPilot-Bench simulator builds upon HighwayEnv [leurent_environment_2018], a widely used platform for autonomous driving research and tactical decision-making, which offers various driving models and simulates realistic multi-vehicle interactions. We extend HighwayEnv with interfaces suitable for LLM-based agents and implement custom intersections to diversify the driving scenarios.
III-A2 Evaluator
The LaMPilot-Bench evaluator incorporates metrics to assess the safety and efficiency of the agent driving policies. Time-to-collision (TTC) measures the vehicle’s ability to maintain safe distances and avoid collisions. In a scenario with $n+1$ vehicles (ego vehicle labeled as 0), the TTC ($\tau_{i}$) with vehicle $i$ is calculated using the ego vehicle’s longitudinal velocity $\mathbf{v}_{0}$ and position $\mathbf{p}_{0}$, and vehicle $i$’s longitudinal velocity $\mathbf{v}_{i}$ and position $\mathbf{p}_{i}$ ($1\leq i\leq n$):
@@ III-A2 Evaluator
Speed variance (SV) is another safety metric. The ego vehicle’s speed standard deviation $\sigma_{0}$ is calculated as:
@@
A TE score closer to 100 indicates a more efficient policy. A task is considered successfully completed when the agent achieves the specified objectives while maintaining safety (i.e., avoiding collisions) and efficiency (i.e., finishing within the allotted time). The final score aggregates all individual metrics, weighted according to their importance:
@@
III-B CARLA Leaderboard 1.0 Benchmark
In addition to LaMPilot-Bench and HighwayEnv, we also utilize the CARLA Leaderboard 1.0 benchmark [carla_team_autonomous_2020, dosovitskiy_carla_2017] to evaluate LLM-based agents in a more realistic and complex driving environment. The Leaderboard consists of 76 routes (50 training, 26 testing) totaling over 170 km across six towns. It also includes predefined scenarios covering ten types of challenging traffic situations, such as control loss and traffic negotiation. To test the instruction-following capabilities of LLM-powered systems, we modify the Leaderboard setup to provide natural language navigation instructions based on the agent’s current position instead of GPS coordinates. To eliminate perception influence on our results, we craft the driving context information using a language generator that leverages CARLA’s privileged information. Executing an LMP in CARLA involves using Python’s exec function, which takes the LMP code as an input string and two dictionaries defining the execution scope: (i) apis, containing all driving APIs the code may call, and (ii) local_vars, initially empty but later holding a generator variable named policy once executed. The Leaderboard evaluates agent performance using three key metrics:
  • Route Completion (RC): measures the percentage of the route distance completed by the agent.
  • Infraction Penalty (IP): tracks various infractions (e.g., collisions, running red lights) committed by the agent, aggregated as a geometric series starting from an ideal base score of 1.0 that is reduced with each infraction.
  • Driving Score (DS): the product of RC and IP, serving as the principal evaluation metric.
III-C Baselines and Results
We establish several baselines to evaluate the performance of LLM-based agents across the benchmarks, as presented in Tab. III for LaMPilot-Bench and Tab. IV for the CARLA Leaderboard 1.0.
@@ III-C Baselines and Results
TABLE IV: Driving performance of expert drivers on CARLA leaderboard 1.0 testing routes. $\uparrow$: Higher values are better.
Method Learning Navigation DS ()($\uparrow$)( ↑ ) RC ()($\uparrow$)( ↑ ) IP ()($\uparrow$)( ↑ )
Roach Expert [zhang_end--end_2021] Reinforcement Learning Waypoint 63.4 100.0 0.63
TCP Expert [wu_trajectory-guided_2022] Reinforcement Learning Waypoint 68.1 95.8 0.71
LLM Planner 0-Shot CoT Instruction 0.00 48.7 0.00
LLM Planner 3-Shot CoT Instruction 17.3 65.3 0.31
LLM Planner Human-Feedback CoT Instruction 65.4 95.1 0.68
Heuristic baselines
@@
Rule-based methods such as the Intelligent Driver Model (IDM)[treiber_congested_2000] and Minimizing Overall Braking Induced by Lane Changes (MOBIL)[kesting_general_2007]. Both methods achieve a zero collision rate in LaMPilot-Bench, as shown in Tab. III. However, it is important to note that these methods operate independently of the provided instructions. Without considering human instructions, IDM and MOBIL achieve success rates of 20.4% and 15.3%, respectively. These results provide a reference point for assessing the effectiveness of LLM-based agents in following human instructions.
Zero-shot and few-shot baselines
LLMs are prompted with task instructions and API documentation, with or without in-context examples. For the few-shot setting, human programmers proficient in Python create in-context examples. They are given API documentation and allowed to write and test their code on the training routes. The same examples are used across all test routes. In the zero-shot setting, GPT models [openai_gpt-4_2023] and PaLM 2 [anil_palm_2023], given only the API Docs, driving context, and instructions, achieve an obvious performance advantage compared to rule-based methods in terms of task completion, as shown in Tab. III. When provided with training examples that include exemplar code (three-shot setting), all the evaluated LLMs exhibit notable improvements in completion rates. For instance, Llama 2 [touvron_llama_2023]’s completion rate increases from 20.4% in the zero-shot setting to 63.3% in the three-shot setting.
Human feedback baselines